A 6.0-cm-diameter horizontal pipe gradually narrows to 4.0 cm . When water flows through this pipe at a certain rate, the gauge pressure in these two sections is 33.0 kPa and 24.0 kPa , respectively. What is the volume rate of flow
Answer #1

from Bernoulli's Principle

P1+(1/2)ρv12+ρgh1=P2+(1/2)ρv22+ρgh2

Since the Pipe is horizontal, h1=h2

So, P1+(1/2)ρv12=P2+(1/2)ρv22

by using continuity equation

A1⋅v1=A2⋅v2
Where v1 and v2 are the velocities of water in the pipe.

Where A1,A2 are the cross sectional areas of the pipe